OpenStudy (i_love_my_nieces):

Help! The solpe of the line that goes though (-3,-5) and(-3,-6) is 1. Positive 2. Negitive 3. Zero 4. Undefined

OpenStudy (luigi0210):

Take the x-coordinates subtract them.. what do you get?

OpenStudy (anonymous):

4

OpenStudy (i_love_my_nieces):

The x- coordinates are the -3's right?

OpenStudy (anonymous):

unidentified

OpenStudy (luigi0210):

\[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] And yea, the -3's. @cgoel Stop giving out direct answers.

OpenStudy (anonymous):

luigi plz help me

OpenStudy (solomonzelman):

cgoel, you are giving out the answers, and the wrong ones !

OpenStudy (i_love_my_nieces):

\[-3 - -3 = 0\]

OpenStudy (solomonzelman):

Yes

OpenStudy (luigi0210):

Right, and what can't you have in the denominator?

OpenStudy (i_love_my_nieces):

Huh?

OpenStudy (solomonzelman):

your denominator in a slope formula is equal to what ?

OpenStudy (luigi0210):

There is a restriction on the denominators.

OpenStudy (anonymous):

could you guys come to my question after u finish up with this one?

OpenStudy (i_love_my_nieces):

I don't know how to find that. @Luigi0210

OpenStudy (luigi0210):

It can't equal 0.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf m= \frac{y_1-y_2}{x_1-x_2} }\) \(\Huge\color{blue}{ \bf m= \frac{(-5)-(-6)}{(-3)-(-3)} }\) also \(\LARGE\color{blue}{ \bf \frac{x}{0}=undefined }\) for any value of x.

OpenStudy (i_love_my_nieces):

I think my answer is Undefined.

OpenStudy (solomonzelman):

Yes, you are thinking correctly. And as I see, you already know why :)

OpenStudy (i_love_my_nieces):

Yes sir. Thank you.

OpenStudy (solomonzelman):

using the slope formula, you can always prove that slope of horizontal line = 0 slope of vertical line = undefined (in other words, it has no slope) I'm not a sir... but..... YOU WELCOME !